OpenStudy (anonymous):

Express the following in simplest radical form. All variables represent positive real numbers. 5/7 square24xy^5 Please break down the steps for me.

OpenStudy (anonymous):

\[\frac{ 5 }{ 7 }\sqrt{24xy^5}\]

OpenStudy (anonymous):

the \(\frac{5}{7}\) out front you can ignore until the end you need to worry about the \(24\) inside the radical, and also the \(y^5\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\sqrt{24}\] is not in simplest radical form because \(24=4\times 6\) making \[\sqrt{24}=\sqrt{4\times 6}=\sqrt{4}\times \sqrt{6}=2\sqrt{6}\] you don't need to write all those steps, that was just by explanation

OpenStudy (anonymous):

\[\sqrt{y^5}\] is not in simplest radical form because the power is 5, which is larger than the index which is 2 (square root) so \[\sqrt{y^5}=\sqrt{y^2\times y^2\times y}=\sqrt{y^2}\times \sqrt{y^2}\times \sqrt{y}\] \[=y\times y\times \sqrt{y}=y^2\sqrt{y}\] again all steps are not necessary, just for explanation

OpenStudy (anonymous):

the snap way to do that one is to say that two goes in to five twice, with a remainder of 1, so that \[\sqrt{y^5}=y^2\sqrt{y}\] i.e. two \(y\)'s come out and one stays in

OpenStudy (anonymous):

wow, this problem is hard but thank you for explaning it to me

OpenStudy (anonymous):

so a "final answer" is \[\frac{5}{7}\times 2\times y^2\sqrt{xy}\] or \[\frac{10y^2}{7}\sqrt{xy}\]

OpenStudy (anonymous):

no i made a mistake, final answer is \[\frac{10y^2}{7}\sqrt{6xy}\]

OpenStudy (anonymous):

it is not that hard with a little practice for example \[\sqrt{x^9}\] is \[x^4\sqrt{x}\] because 2 goes in to 9 4 times, with a remainder of 1 so 4 come out, one stays in

OpenStudy (anonymous):

oh okay, i kind of getting it

OpenStudy (anonymous):

works with larger indexes as well \[\sqrt[3]{x^{11}}=x^3\sqrt{x^2}\]because 3 goes in to 11 3 times with a remainder of 2, so three come out, two stay in

OpenStudy (anonymous):

also \[\sqrt{50}=5\sqrt{2}\] because \(50=25\times 2\) and the square root of 25 you know is 5

OpenStudy (anonymous):

Again, thanks